# Preprint Notice

### v1

@@ Abstract
The pursuit of high-performance data transfer often focuses on raw network bandwidth, and international links of 100 Gbps or higher are frequently considered the primary enabler. While necessary, this network-centric view is incomplete, equating provisioned link speeds with practical, sustainable data movement capabilities across the entire edge-to-core spectrum. This paper investigates six common paradigms, from the often-cited constraints of network latency and TCP congestion control algorithms to host-side factors such as CPU performance and virtualization that critically impact data movement workflows. We validated our findings using a latency-emulation-capable testbed for high-speed WAN performance prediction and through extensive production measurements from resource-constrained edge environments to a 100 Gbps operational link connecting Switzerland and California, U.S. These results show that the principal bottlenecks often reside outside the network core, and that a holistic hardware-software co-design ensures consistent performance, whether moving data at 1 Gbps or 100 Gbps and faster. This approach effectively closes the fidelity gap between benchmark results and diverse and complex production environments.
I Introduction:
@@
The efficient movement of data from edge locations to core data centers and cloud resources is a foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for their raw bandwidth, are a critical response to this challenge. For example, ESnet6 (the 6th generation of the Energy Sciences Network backbone, operated by the U.S. Department of Energy) provides high-capacity transcontinental and international connectivity essential for global data-intensive scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps [1]. However, a predominant focus on raw network bandwidth can obscure a more critical truth: sustainable end-to-end data movement is constrained by the full environment encountered along the data path—including storage, host architectures, software design, and implemented security measures—rather than by the network alone.
Drawing on over a decade of operational experience—spanning extreme scientific workloads at Linac Coherent Light Source II (LCLS-II) [2], sustained petabyte transfers [3] and [4], international deployments [5], and resource-constrained edge environments, all with consistent near-line-rate performance—our work demonstrates that once a “fast” or “high-speed” link is in place, storage subsystems and host configurations become the primary determinants of performance for file-based transfers. This is evidenced by achieving near-line-rate performance on a “high-speed” 100 Gbps transcontinental link (~84 Gbps in a pharmaceutical production environment) while also enabling high-efficiency transfers at 1-10 Gbps rates at the edge. In this paper, we define a 'fast' network as 10 Gbps up to (but excluding) 100 Gbps, and a 'high-speed' network as 100 Gbps and faster.
@@
A key architectural component for achieving predictable performance is the use of high-speed burst buffers [6] as staging areas, enabling reliable, near-line-rate movement when the end-to-end path is properly co-designed. Our prior work, including the Overall Winner at the Inaugural Supercomputing Asia 2019 Data Mover Challenge (SCA19 DMC) [7] and ESnet official evaluation [8], underscores that these factors, not the network, set the true performance limits of data movement, regardless of the network's nominal speed.
@@
The core contribution of this paper is to identify and empirically validate a set of widely held engineering beliefs—six specific paradigms—that continue to shape expectations and investments in production end-to-end data movement. This work examines why these paradigms fall short across the full performance spectrum. We demonstrate that the path to reliable performance lies not in isolated component optimization, but in holistic, co-designed and highly-integrated hardware and software systems.
@@
The six paradigms we re-examine are: (1) the critical impact of network latency, (2) the prevalence and impact of packet loss and the criticality of TCP Congestion Control Algorithms (CCA), (3) the necessity of dedicated private lines for high-speed testing, (4) the direct relationship between increased network bandwidth and transfer rates, (5) the essential need for powerful CPUs, and (6) the universal utility of virtualization and cloud environments. Throughout this paper, all referenced systems are assumed to run Red Hat Enterprise Linux (RHEL) 9.6 [9] or free rebuilds such as Rocky Linux 9.6 [10].
@@ III-A The Latency Killer paradigm
Figure 1: iperf3 latency sweep results obtained using two HPE DL380 Gen 11 server-based appliances (Fig. 2) on a latency simulation-capable testbed (max network speed 100 Gbps) established in the former Intel Swindon Lab, Swindon, U.K., in 2024. While default kernel networking settings (OOTB) show severe performance degradation under high latency, proper kernel tuning substantially reduces this penalty.”
@@ III-F1 The Architectural Cost of Cloud Abstraction
Standard cloud workarounds, such as AWS multipart uploads [49], fragment a simple file transfer into a multi-step coordination process—effectively a Rube Goldberg machine [50] for data movement.
@@ III-F2 A Critical Examination of Cloud Provider Metrics
The prevailing paradigm of general-purpose cloud data paths, regardless of provider, introduces inherent architectural conflicts for high-performance data movement workflows. This challenge becomes acute at the petascale level. Analysis of major U.S. cloud providers reveals a focus on theoretical network bandwidth, often decoupled from storage I/O and end-to-end performance. The info in Tables VII-IX, extracted from official provider documentation, illustrate this focus.
@@
TABLE VIII: Amazon Web Services (AWS) virtual instance parameters.
@@ III-F4 Empirical Evidence of the Cloud Data Movement Limitations
The performance penalties imposed by standard cloud data paths are measurable. Using the GoToCloud platform on AWS [53], a comparative analysis conducted by High Energy Accelerator Research Organization (KEK) [54] vividly illustrates this point. As summarized in Fig. 11, the transfer of a 1.2 TiB Cryo-EM dataset was measured with zx vs. aws-cli
@@ IV Methodology and Reproducibility
The Zettar zx data mover and associated deployment tooling represent proprietary commercial technology reserved for customers and partners. The provided materials allow independent verification of our benchmarking methodology without requiring access to our commercial software.
@@ V Acknowledgment
  • Dr. Jana Thayer (LCLS Experimental Data Systems Division Director) for her review and for insights into the LCLS-II data system architecture and real-time processing requirements that motivated our focus on streaming data movement and holistic system design.
  • Dr. Wilko Kroeger (LCLS-II Information Specialist, SLAC) for his insights and review of the data management challenges associated with the Linac Coherent Light Source II (LCLS-II), which informed the architectural perspectives in this work.

### v2

@@ Abstract
The pursuit of high-performance data transfer often focuses on raw network bandwidth, where international links of 100 Gbps or higher are frequently considered the primary enabler. While necessary, this network-centric view is incomplete, as it equates provisioned link speeds with practical, sustainable data movement capabilities across the entire edge-to-core spectrum. This paper investigates six common paradigms, ranging from network latency and TCP congestion control to host-side factors such as CPU performance and virtualization that critically impact data movement workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. We introduce the “Drainage Basin Pattern” conceptual model for reasoning about end-to-end data flow constraints across heterogeneous hardware and software components to address the fidelity gap between raw bandwidth and application-level throughput. Our findings are validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations and transcontinental production trials over 100 Gbps operational links. The results demonstrate that principal bottlenecks often reside outside the network core, and that a holistic hardware-software co-design enables consistent, predictable performance for moving data at scale and speed.
I Introduction
@@
The efficient movement of data from edge locations to core data centers and cloud resources is a foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for their raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps [1]. However, a predominant focus on raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path—including storage, host architectures, software design, and security measures—rather than by the network alone.
We identify this discrepancy as a “fidelity gap” between theoretical link capacity and actual application-level throughput. This gap is addressed by the “Drainage Basin Pattern” conceptual model (Fig. 1), which shifts the focus from isolated network optimization to systemic co-design. The efficacy of this approach is evidenced by achieving near-line-rate performance on a “high-speed” 100 Gbps transcontinental link (~84 Gbps in a pharmaceutical production environment) while enabling high-efficiency transfers at 1–10 Gbps at the edge. In this paper, we define a “fast” network as 10 Gbps up to (but excluding) 100 Gbps, and a “high-speed” network as 100 Gbps and faster.
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II [2], sustained petabyte transfers [3, 4], and international deployments [5]—our work demonstrates that once a high-speed link is in place, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations [8, 14, 15] and historical transcontinental trials over 100 Gbps operational links.
@@
A key architectural component for achieving predictable performance is the use of high-speed burst buffers [6] as staging areas, enabling reliable, near-line-rate movement when the end-to-end path is properly co-designed. Our prior work, including the Overall Winner at the Inaugural Supercomputing Asia 2019 Data Mover Challenge (SCA19 DMC) [7] and official ESnet evaluations [8, 14, 15], underscores that these factors, not the network, set the true performance limits of data movement, regardless of the network’s nominal speed.
@@
The core contribution of this paper is to identify and validate a set of widely held engineering beliefs—six specific paradigms—that continue to shape expectations and investments in production end-to-end data movement. This work examines why these paradigms often fall short across the full performance spectrum. We demonstrate that the path to reliable and predictable performance lies not in isolated component optimization, but in holistic, co-designed, and highly-integrated hardware and software systems.
@@
The six paradigms we re-examine are: (1) the assumed dominance of network latency as the primary performance constraint; (2) the impact of packet loss and the perceived sufficiency of standard TCP Congestion Control Algorithms (CCA); (3) the presumed necessity of dedicated private lines for high-speed performance verification; (4) the linear expectation of a direct relationship between increased link bandwidth and application transfer rates; (5) the belief in the essential need for high-performance CPUs at every endpoint; and (6) the assumed universal utility of virtualization and cloud environments for high-throughput workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. Throughout this paper, all referenced systems are assumed to run Red Hat Enterprise Linux (RHEL) 9.6 [9] or free rebuilds such as Rocky Linux 9.6 [10].
@@ III-A The Latency Killer Paradigm
Figure 2: iperf3 latency sweep results obtained using two HPE DL380 Gen 11 server-based appliances (Fig. LABEL:fig:fig) on a latency simulation-capable testbed (max network speed 100 Gbps) established in the former Intel Swindon Lab, Swindon, U.K., in 2024. While default kernel networking settings (OOTB) show severe performance degradation under high latency, proper kernel tuning substantially reduces this penalty.”
@@ III-F1 The Architectural Cost of Cloud Abstraction
Standard cloud workarounds, such as Amazon Web Services (AWS) multipart uploads [49], fragment a simple file transfer into a multi-step coordination process—effectively a Rube Goldberg machine [50] for data movement.
@@ III-F2 A Critical Examination of Cloud Provider Metrics
The prevailing paradigm of general-purpose cloud data paths, regardless of provider, introduces inherent architectural conflicts for high-performance data movement workflows. This challenge becomes acute at the petascale level. Analysis of major U.S. cloud providers reveals a focus on theoretical network bandwidth, often decoupled from storage I/O and end-to-end performance. The info in Tables VII-IX, extracted from official provider documentation, illustrates this focus.
@@
TABLE VIII: AWS virtual instance parameters.
@@ III-F4 Empirical Evidence of the Cloud Data Movement Limitations
The performance penalties imposed by standard cloud data paths are measurable. Using the GoToCloud platform on AWS [53], a comparative analysis conducted by High Energy Accelerator Research Organization (KEK) [54] vividly illustrates this point. As summarized in Fig. 11, the transfer of a 1.2 TiB Cryo-electron microscopy (Cryo-EM) dataset was measured with zx vs. aws-cli
@@ IV Methodology and Reproducibility
The zx data mover and associated deployment tooling represent proprietary commercial technology reserved for customers and partners. The provided materials allow independent verification of our benchmarking methodology without requiring access to our commercial software.
@@ V Acknowledgment
  • Dr. Jana Thayer (LCLS Experimental Data Systems Division Director, SLAC) for her review and for insights into the LCLS-II data system architecture and real-time processing requirements that motivated our focus on streaming data movement and holistic system design.
  • Dr. Wilko Kroeger (LCLS-II Information Specialist, SLAC) for his insights and review of the data management challenges associated with the LCLS-II, which informed the architectural perspectives in this work.